### v1

# BOLIMES: Boruta–LIME optiMized fEature Selection for Gene Expression Classification

Abstract
Gene expression classification is a pivotal yet challenging task in bioinformatics, primarily due to the high dimensionality of genomic data and the risk of overfitting. To bridge this gap, we propose BOLIMES, a novel feature selection algorithm designed to enhance gene expression classification by systematically refining the feature subset. Unlike conventional methods that rely solely on statistical ranking or classifier-specific selection, we integrate the robustness of Boruta with the interpretability of LIME, ensuring that only the most relevant and influential genes are retained. BOLIMES first employs Boruta to filter out non-informative genes by comparing each feature against its randomized counterpart, thus preserving valuable information. It then uses LIME to rank the remaining genes based on their local importance to the classifier. Finally, an iterative classification evaluation determines the optimal feature subset by selecting the number of genes that maximizes predictive accuracy. By combining exhaustive feature selection with interpretability-driven refinement, our solution effectively balances dimensionality reduction with high classification performance, offering a powerful solution for high-dimensional gene expression analysis.
Keywords:
Image Classification Gene Expression Boruta LIME Feature Selection
1 Introduction
Gene expression classification [1, 17, 9, 10, 18] has emerged as a fundamental tool in bioinformatics, enabling the identification of disease subtypes, prediction of patient outcomes, and discovery of potential therapeutic targets. With the advent of high-throughput sequencing technologies, researchers can now analyze vast gene expression profiles across thousands of genes simultaneously. However, this progress comes with a significant computational and analytical challenge: the curse of dimensionality [22, 2]. In typical gene expression datasets, the number of genes ($p$) far exceeds the number of samples ($n$), often by several orders of magnitude. This imbalance leads to severe overfitting in machine learning models, where classifiers struggle to generalize due to the overwhelming presence of irrelevant or redundant features. Additionally, the high dimensionality increases computational complexity, making conventional classification models inefficient and impractical for real-world applications.
To address these challenges, feature selection plays a critical role in gene expression classification. By identifying and retaining only the most informative genes, feature selection not only improves model generalization but also enhances interpretability, allowing researchers to derive biologically meaningful insights from machine learning predictions [9, 20]. Furthermore, reducing the feature space significantly lowers computational costs, enabling faster model training and inference. More importantly, gene selection aligns with the biological reality that only a subset of genes actively contributes to disease mechanisms, making feature selection a crucial step in biomedical analysis. Without an effective selection strategy, classifiers are prone to noise, reduced accuracy, and difficulty in extracting meaningful biomarkers for clinical applications.
Given the importance of feature selection, extensive research has been conducted to develop robust selection techniques tailored for gene expression data. Traditional methods include filter-based approaches [11, 5, 8, 26], which rely on statistical measures such as mutual information, correlation, and entropy to rank features independently of the classifier. Although computationally efficient, filter methods often fail to capture complex gene interactions. Wrapper-based methods [3, 33, 21, 28], such as recursive feature elimination (RFE) [7, 34] and genetic algorithms [25], iteratively refine feature subsets based on classifier performance but tend to be computationally expensive for high-dimensional data. Meanwhile, embedded techniques [19, 31, 16], such as LASSO [13, 29] and tree-based models [14, 27], integrate feature selection within the classification process but may not always provide optimal feature subsets for different classifiers. Despite these advancements, no single approach consistently outperforms others across all gene expression datasets, necessitating hybrid strategies that leverage the strengths of multiple selection paradigms.
To bridge this gap, we propose BOLIMES (Boruta–LIME Enhanced eXtraction), a novel feature selection algorithm designed to enhance gene expression classification by systematically refining the feature subset. Unlike conventional methods that rely solely on statistical ranking or classifier-specific selection, BOLIMES integrates the robustness of Boruta [24, 36] with the interpretability of LIME [30], ensuring that only the most relevant and influential genes are retained. BOLIMES first applies Boruta, a powerful wrapper method that iteratively refines feature sets by comparing real genes against randomly permuted shadow features, ensuring that no informative gene is prematurely discarded. Then, LIME [30, 23] (Local Interpretable Model-agnostic Explanations) is employed to assess the local importance of each selected gene, providing an interpretable ranking of features that capture model-specific contributions. Finally, an iterative classification evaluation determines the optimal feature subset by selecting the number of genes that maximize predictive accuracy. By integrating Boruta’s exhaustive feature selection with LIME’s interpretability-driven refinement, BOLIMES effectively balances dimensionality reduction with model performance, making it a powerful solution for high-dimensional gene expression analysis.
The remainder of this paper is structured as follows: we briefly presents a fundamental background in Section 2. Next, Section 3 describes our proposal algorithm. Then, Section 4 represents the experiment, the results of the models and an discussion. Finally, Section 5 shows the conclusions and future work.
2 Background
In this section, we present the foundation of our approach with integrating Boruta and Lime. We also provide the ML algorithm that will be used for the gene expression classification task.
2.1 Gene Expression Classification with ML models
Gene expression classification [9, 10, 18] lies at the forefront of biomedical research, offering profound insights into the molecular mechanisms underlying various diseases. ML models have become indispensable in this domain, as they can uncover complex patterns within vast and high-dimensional gene expression datasets. However, these datasets often contain a plethora of features, many of which are redundant or irrelevant, potentially obscuring the most critical biological signals and leading to overfitting. Consequently, feature selection becomes imperative—it refines the dataset by isolating the most informative genes, thereby enhancing model accuracy, interpretability, and computational efficiency. By focusing solely on the pivotal biomarkers, this research is able to achieve more reliable predictive outcomes. In this paper, we investigate and evaluate the classification with various ML techniques. Namely, we experiment our selected features with ML algorithms, i.e., SVM [32], Random Forest [4], XGB [6], Gradient Boosting [15].
Definition 1 (Classification)
Let $D=(X,y)$ be a dataset where $X\subseteq\mathbb{R}^{n}$ is the feature space and $y\in\mathcal{Y}=\{1,2,\dots,k\}$ represents the class labels. A classifier is a function
$f:X\to\mathcal{Y},$
that assigns a predicted label $\hat{y}=f(x)$ to each input $x\in X$. The function $f$ is learned from the labeled examples
$D=\{(x_{i},y_{i})\mid x_{i}\in X,\;y_{i}\in\mathcal{Y},\;i=1,\dots,N\},$
by minimizing a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ that quantifies the error between the predicted and true labels. Once trained, $f$ is used to classify new, unseen inputs.
While numerous feature selection techniques exist, our study concentrates on two: Boruta and LIME. We choose Boruta for its robust ability to identify all truly relevant features within high-dimensional datasets, ensuring that no significant predictor is overlooked. Complementing this, LIME is employed for its exceptional capacity to provide local, interpretable explanations of model predictions, which is crucial when assessing feature importance through numerous distance calculations. Together, these methods offer a powerful and efficient approach to refining our gene expression classification models. Now, we introduce Boruta and Lime in the next session.
2.2 Leveraging Boruta for Robust Feature Extraction
Boruta [24, 36] is a powerful wrapper-based feature selection algorithm designed to identify all truly relevant variables in a dataset. By comparing the importance of actual features with that of randomly generated “shadow” features, Boruta systematically filters out irrelevant variables while preserving essential predictors. This rigorous selection process is particularly valuable in high-dimensional applications, such as gene expression classification, where capturing meaningful signals is crucial. Rather than directly improving predictive accuracy, Boruta refines the feature set, which can indirectly enhance model interpretability and performance. For clarity, we formally define Boruta as follows:
Definition 2 (Boruta Feature Selection)
Let $D=(X,y)$ be a dataset with features $X=\{x_{1},x_{2},\dots,x_{p}\}$ and target $y$. The Boruta algorithm identifies all relevant features in $X$ as follows:
  1. 1.
    Shadow Feature Generation: For each $x_{i}\in X$, create a shadow feature $x_{i}^{\text{shadow}}$ by randomly permuting its values, forming the set $X^{\text{shadow}}$.
  2. 2.
    Importance Estimation: Train a classifier (e.g., Random Forest) on the combined set $X\cup X^{\text{shadow}}$ and compute the importance score $I(z)$ for each $z$.
  3. 3.
    Feature Comparison: For each $x_{i}$, define
    $I^{\text{shadow}}_{\max}=\max_{z\in X^{\text{shadow}}}I(z).$
    Then classify $x_{i}$ as relevant if $I(x_{i})$ is significantly greater than $I^{\text{shadow}}_{\max}$, irrelevant if significantly lower, or tentative otherwise.
  4. 4.
    Iteration: Remove irrelevant features and repeat until all features are decisively classified.
The final selected subset $X^{*}\subseteq X$ comprises all features deemed relevant.
2.3 XAI for Feature Selection
Explainable AI (XAI) [12, 35] represents a forefront of AI research, aiming to elucidate the decision-making processes of complex models. In the context of gene expression classification, where feature selection is pivotal to model performance and interpretability, our study leverages LIME—Local Interpretable Model-Agnostic Explanations—to demystify and extract critical features. LIME approximates the behavior of a sophisticated, black-box model with a simpler, locally interpretable surrogate, thereby pinpointing the most influential predictors in the vicinity of a given instance. This approach enhances the transparency of the model’s predictions andfacilitates a more informed and rigorous feature selection process, ultimately contributing to both improved accuracy and trustworthiness of the classification system. Now, we provide a formal definition of LIME as follows:
Definition 3 (LIME-based Feature Selection)
Let $D^{*}=(X^{*},y)$ be the dataset resulting from Boruta, where $X^{*}\subseteq\mathbb{R}^{p^{*}}$ is the set of relevant features. Given a trained black-box classifier $f:X^{*}\to\mathcal{Y}$ and an instance $x\in X^{*}$, LIME constructs an interpretable surrogate model $g\in G$ (typically linear), expressed as
$g(z)=w_{0}+\sum_{j=1}^{p^{*}}w_{j}z_{j},$
by solving the optimization problem
$\min_{g\in G}\sum_{z\in Z_{x}}\pi_{x}(z)\left(f(z)-g(z)\right)^{2}+\Omega(g),$
where $Z_{x}$ is a set of perturbed samples in the neighborhood of $x$, $\pi_{x}(z)$ is a proximity measure, and $\Omega(g)$ enforces simplicity. The absolute coefficients $|w_{j}|$ indicate the local importance of each feature, enabling a further refined selection $X_{\text{selected}}\subseteq X^{*}$ for classification.
3 BOLIMEX algorithm
Our methodology begins with applying the Boruta algorithm to sift through the high-dimensional gene expression dataset, effectively filtering out irrelevant features and isolating those that are truly significant. This initial reduction is critical because LIME, our subsequent interpretability tool, involves generating numerous perturbed samples and calculating distances—a computationally intensive process, especially in large feature spaces. By narrowing the focus to a refined subset via Boruta, we significantly reduce the computational burden and enhance the precision of LIME’s local explanations. This sequential approach streamlines the overall feature selection process and fortifies the reliability and clarity of our model’s interpretability, ultimately leading to improved classification performance.
However, the key question remains: how many features should be selected for optimal classification? In our proposed BOLIMES algorithm, which integrates Boruta and LIME, we first reduce the high-dimensional feature set by eliminating irrelevant variables with Boruta. Then, we further refine this subset using LIME to assess the local importance of each feature. Finally, we determine the optimal number of features by evaluating classification performance—selecting the subset that yields the highest accuracy for model training. In general, our model is both efficient and robust, relying only on the most informative features for gene expression classification.
1
Input : Dataset $D=(X,y)$ with $X\subseteq\mathbb{R}^{p}$ and class labels $y$.
Output : Optimal feature subset $X_{\text{opt}}$ and trained classifier $f_{\text{opt}}$.
2
3begin
       $X^{*}\leftarrow\text{Boruta}(D)$
       // Identify relevant features from $X$
4
      $\mathcal{I}\leftarrow\text{LIME}(f,X^{*})$
       // Compute local importance scores on $X^{*}$
5
      $X^{*}_{R}\leftarrow\{x^{*}_{(i)}\}_{i=1}^{|X^{*}|}$$\mathcal{I}(x^{*}_{(1)})\geq\dots\geq\mathcal{I}(x^{*}_{(|X^{*}|)})$
       // Rank features in $X^{*}$ in descending order of $\mathcal{I}$
6
[…]
8
9      for $k=1$ to $|X^{*}|$ do
             $S_{k}\leftarrow\{x^{*}_{(i)}\}_{i=1}^{k}$
             // Select top-$k$ features from $X^{*}_{R}$
             $f_{k}\leftarrow\text{TrainClassifier}(D_{S_{k}})$
             // Train classifier on $D_{S_{k}}$
             $\text{acc}\leftarrow\text{Evaluate}(f_{k},D_{S_{k}})$
             // Compute classification accuracy
10
[…]
13                   $k^{*}\leftarrow k$
14                   $f_{\text{opt}}\leftarrow f_{k}$
15
      $X_{\text{opt}}\leftarrow\{x^{*}_{(i)}\}_{i=1}^{k^{*}}$
       // Select top-$k^{*}$ features from $X^{*}_{R}$
16
17      return $X_{\text{opt}},f_{\text{opt}}$
Algorithm 1 Optimal Feature Selection for Classification using Boruta and LIME
The algorithm presents an optimal feature selection framework by integrating Boruta and LIME to refine feature subsets for classification. Given a dataset $D=(X,y)$, it first applies Boruta to extract a subset $X^{*}$ of relevant features. LIME then computes local importance scores $\mathcal{I}$ for $X^{*}$, producing a ranked set $X^{*}_{R}$ where
$X^{*}_{R}\leftarrow\{x^{*}_{(i)}\}_{i=1}^{|X^{*}|},\quad\mathcal{I}(x^{*}_{(1)})\geq\dots\geq\mathcal{I}(x^{*}_{(|X^{*}|)}).$
An iterative search determines the optimal number of features $k^{*}$ by selecting the top-$k$ ranked features, training a classifier $f_{k}$, and evaluating its accuracy:
$S_{k}\leftarrow\{x^{*}_{(i)}\}_{i=1}^{k},\quad f_{k}\leftarrow\text{TrainClassifier}(D_{S_{k}}).$
The best-performing classifier defines $k^{*}$, yielding the final subset $X_{\text{opt}}$ and trained model $f_{\text{opt}}$.
The algorithm’s complexity is driven by three key stages: Boruta for feature selection, LIME for ranking, and iterative classification. Boruta, relying on multiple iterations of Random Forest, has a worst-case complexity of $O(T\cdot p^{2}\log p)$. LIME, which perturbs $m$ samples per feature, contributes $O(m\cdot p)$. The final stage trains classifiers iteratively over $p^{*}$ ranked features, leading to an overhead of $O(np^{*}{}^{2})$ assuming a model with $O(nk)$ complexity. Thus, the total complexity is: $O(T\cdot p^{2}\log p)+O(m\cdot p)+O(np^{*}{}^{2}),$ where $p^{*}\ll p$ in practice, making the approach feasible for moderate-dimensional data but computationally intensive for extremely large $p$.
4 Experiment and results
This section offers a brief description of the gene expression datasets while delivering a detailed comparative analysis of the classification models. Furthermore, we also provide the results of Boruta. Our source code has been made publicly accessible on GitHub [^1] https://github.com/Vannguyen0312/Manholes-App/.
4.1 Dataset and Configurations
The gene expression datasets summarized in Table 1 exemplify the inherent challenges of high-dimensional biomedical data. With sample sizes ranging from 53 to 575 and feature counts spanning from approximately 11,950 to over 54,600, these datasets present a significant imbalance between the number of available samples and the vast dimensionality of gene expression profiles. Additionally, the variability in the number of classes—from as few as 3 to as many as 14—further complicates the classification task by introducing diverse and complex biological signals. This high dimensionality coupled with limited sample sizes accentuates the risk of overfitting and underscores the critical need for effective feature selection. Robust feature selection methods are essential to isolate the most informative genes, thereby enhancing model interpretability and predictive accuracy in gene expression classification. In this study, our feature selection strategy is specifically designed to address these challenges, ensuring that only the most relevant features are retained for subsequent classification tasks.
Table 1: Dataset Characteristics
ID Dataset #Datapoints #Dimensions #Classes
1 E-GEOD-20685 327 54627 6
2 E-GEOD-20711 90 54675 5
3 E-GEOD-21050 310 54613 4
4 E-GEOD-21122 158 22283 7
5 E-GEOD-29354 53 22215 3
6 E-GEOD-30784 229 54675 3
7 E-GEOD-31312 498 54630 3
8 E-GEOD-31552 111 33297 3
9 E-GEOD-32537 217 11950 7
10 E-GEOD-33315 575 22283 10
11 E-GEOD-36895 76 54675 14
12 E-GEOD-37364 94 54675 4
13 E-GEOD-39582 566 54675 6
14 E-GEOD-39716 53 33297 3
15 E-GEOD-44077 226 33252 4
To evaluate our approach and the training model, we implement the AI libraries (i.e., tensorflow (version 3.10.0), Keras (version 3.0)) and run the experiments on the computer with the following configuration: AMD Ryzen 7 6800HS, Ceator Edition 3.20 GHZ, 16 GB RAM, Windows 10 OS.
4.2 Boruta’s Feature Selection
In this study, we adopt Boruta as our primary feature selection method to substantially reduce the dimensionality of our gene expression datasets before applying LIME for model interpretation. The rationale behind this choice is to prevent an explosion in computational complexity and potential loss of interpretability when LIME is applied to an excessively high-dimensional feature space. By using Boruta, we are able to effectively eliminate irrelevant features while retaining those that are truly informative for classification. As evidenced in Table 2, datasets such as E-GEOD-20685 were reduced from 54,627 dimensions to only 545 confirmed features, thereby rendering the subsequent LIME analysis both feasible and efficient. To achieve this, Boruta was executed with the following parameters via the BorutaPy library: a random forest classifier (‘rf‘) with 300 estimators, a maximum of 200 iterations, an alpha value of ’0.01’, a percentile threshold of 95, two-step feature selection enabled, and a fixed random state of 42, with verbose output enabled. These parameter settings were meticulously chosen to ensure a rigorous and robust selection process, ultimately facilitating a more interpretable and high-performing classification model.
Table 2: Feature Selection Results
ID Dataset Features Feature Selection
E-GEOD-* Confirmed Tentative Rejected Time (s)
1 20685 545 262 53820 633,0205021
2 20711 111 33 54531 140,0120337
3 21050 72 17 54524 350,3323104
4 21122 271 124 21888 219,2371044
5 29354 28 3 22184 91,59759808
6 30784 171 64 54440 279,9282582
7 31312 213 42 54375 777,7057292
8 31552 79 48 33170 136,9728532
9 32537 96 17 11837 232,6542134
10 33315 483 108 21692 1176,811348
11 36895 39 14 54622 122,578902
12 37364 59 43 54573 129,0980539
13 39582 640 253 53782 1306,720685
14 39716 124 29 33144 99,24336505
15 44077 227 231 32794 221,3678348
4.3 Classification results
In our study, we conducted extensive experiments with four advanced algorithms to determine the method yielding the highest accuracy across various gene expression datasets. We configured an SVM using SVC with a radial basis function kernel, setting C to 100,000, gamma to ”scale”, and class weights to ”balanced”, ensuring reproducibility with a random state of 42. In parallel, we employed a Random Forest classifier with 200 estimators, a maximum depth of 10, and balanced class weights. Additionally, we implemented both an XGBClassifier and a GradientBoostingClassifier, each configured with 50 estimators, a maximum depth of 10, and a learning rate of 0.01, with the random state fixed at 42. Our comparative evaluation revealed that while the SVM excelled on datasets with fewer samples, the ensemble methods—particularly Random Forest and XGBoost—demonstrated more robust performance on high-dimensional data. This comprehensive analysis highlights the critical role of algorithm selection and fine-tuning in achieving optimal classification performance in gene expression studies.
Table 3: SVM Classification Results with Top K Features
ID Dataset Class Method Samples Top K Classification Results Time (s)
Features Acc Prec Rec F1 Score Training Preprocess Total
ours 327 107 0,955 0,956 0,955 0,955 1,151 12,951 112,137
1 20685 6 327 107 0,955 0,956 0,955 0,955 1,151 12,951 112,137
327 107 0,955 0,956 0,955
2 20711 5 90 67 0,944 0,952 0,944 0,944 0,298 3,745 17,754
3 21050 4 310 70 0,677 0,710 0,677 0,678 1,494 10,400 295,982
4 21122 7 158 170 0,938 0,915 0,938 0,922 0,548 2,010 82,092
5 29354 3 53 10 1,000 1,000 1,000 1,000 0,165 0,823 0,165
6 30784 3 229 14 1,000 1,000 1,000 1,000 0,730 8,197 3,476
7 31312 3 498 43 0,900 0,900 0,900 0,900 0,063 19,236 2,144
8 31552 3 111 15 0,957 0,917 0,957 0,936 0,299 2,642 1,816
9 32537 7 217 40 0,796 0,730 0,796 0,747 0,735 0,918 21,201
10 33315 10 575 273 0,878 0,860 0,878 0,867 0,138 4,844 33,250
11 36895 14 76 15 0,929 0,893 0,929 0,905 0,228 3,199 1,304
12 37364 4 94 17 0,947 0,965 0,947 0,947 0,259 3,393 2,083
13 39582 6 566 251 0,877 0,889 0,877 0,880 0,130 20,512 25,058
14 39716 3 53 12 1,000 1,000 1,000 1,000 0,157 1,642 0,454
15 44077 4 226 18 1,000 1,000 1,000 1,000 0,578 5,513 5,220
The classification results obtained using SVM, Random Forest, XGBoost, and Gradient Boosting reveal distinct performance and computational trade-offs. SVM generally delivers high accuracy and balanced precision, recall, and F1 scores across most datasets, with relatively low training times in some cases; however, its preprocessing time can be substantial when handling a large number of features. In contrast, Random Forest exhibits competitive accuracy, often outperforming SVM on certain datasets, yet it tends to incur significantly higher training times, particularly when the number of top features increases—as observed in dataset 1 where training time was notably high. XGBoost demonstrates robust performance with competitive accuracy and precision; its moderate training and preprocessing times suggest that it strikes a balance between model complexity and computational efficiency, though its performance can be dataset-dependent (e.g., lower accuracy on dataset 13). Meanwhile, Gradient Boosting achieves near-perfect accuracy on several datasets, indicating high robustness in classification, yet its training and total processing times vary considerably across datasets.
Table 4: Random Forest Classification Results with Top K Features
ID Dataset Class Samples Top N Classification Results Time (s)
Features Acc Prec Rec F1 Score Training Preprocess Total
1 20685 6 327 487 0,970 0,971 0,970 0,970 40,762 13,055 16306,229
2 20711 5 90 10 0,833 0,841 0,833 0,830 5,396 3,703 5,396
3 21050 4 310 72 0,726 0,727 0,726 0,711 25,093 10,252 1507,624
4 21122 7 158 104 0,938 0,922 0,938 0,925 11,577 1,971 1039,209
5 29354 3 53 10 1,000 1,000 1,000 1,000 3,325 1,059 3,325
6 30784 3 229 12 1,000 1,000 1,000 1,000 14,531 8,278 43,917
7 31312 3 498 161 0,900 0,821 0,900 0,858 1,594 19,739 221,322
8 31552 3 111 12 1,000 1,000 1,000 1,000 6,667 2,608 20,318
9 32537 7 217 87 0,773 0,680 0,773 0,707 17,146 0,935 1238,228
10 33315 10 575 444 0,887 0,854 0,887 0,868 3,215 4,947 985,866
11 36895 14 76 20 0,929 0,929 0,929 0,929 4,347 3,242 46,355
12 37364 4 94 10 0,947 0,953 0,947 0,947 5,579 3,380 5,579
13 39582 6 566 189 0,877 0,890 0,877 0,876 1,908 20,970 304,137
14 39716 3 53 15 1,000 1,000 1,000 1,000 3,172 1,528 19,338
15 44077 4 226 52 1,000 1,000 1,000 1,000 15,408 6,125 642,643
Table 5: XGB Classification Results with Top K Features
IID Dataset Class Samples Top N Classification Results Time (s)
Features Acc Prec Rec F1 Score Training Preprocess Total
1 20685 6 327 114 0,955 0,962 0,955 0,955 0,395 14,439 40,015
2 20711 5 90 10 0,833 0,841 0,833 0,830 0,328 3,800 0,328
3 21050 4 310 60 0,790 0,791 0,790 0,784 0,286 11,975 16,186
4 21122 7 158 159 0,938 0,921 0,938 0,925 0,293 4,028 44,087
5 29354 3 53 10 1,000 1,000 1,000 1,000 0,271 2,541 0,271
6 30784 3 229 50 0,978 0,979 0,978 0,976 0,127 10,126 6,344
7 31312 3 498 10 0,860 0,848 0,860 0,843 0,388 21,554 0,388
8 31552 3 111 10 1,000 1,000 1,000 1,000 0,557 4,239 0,557
9 32537 7 217 95 0,796 0,693 0,796 0,736 0,299 2,478 26,388
10 33315 10 575 425 0,870 0,842 0,870 0,855 1,341 6,495 389,225
11 36895 14 76 29 0,929 0,929 0,929 0,929 0,250 3,364 6,509
12 37364 4 94 21 1,000 1,000 1,000 1,000 0,142 4,804 1,862
13 39582 6 566 556 0,825 0,833 0,825 0,824 1,564 22,414 554,762
14 39716 3 53 44 1,000 1,000 1,000 1,000 0,101 3,074 3,746
15 44077 4 226 57 0,978 0,980 0,978 0,979 0,139 6,925 8,332
Table 6: Gradient Boosting Classification Results with Top N Features
IID Dataset Class Samples Top N Classification Results Time (s)
Features Acc Prec Rec F1 Score Training Preprocess Total
1 20685 6 327 468 0,939 0,947 0,939 0,936 17,957 12,795 17,957
2 20711 5 90 38 0,889 0,921 0,889 0,892 0,664 3,832 0,664
3 21050 4 310 65 0,807 0,802 0,807 0,804 1,844 9,923 1,844
4 21122 7 158 257 0,875 0,872 0,875 0,871 5,544 1,890 5,544
5 29354 3 53 12 1,000 1,000 1,000 1,000 0,387 0,822 0,387
6 30784 3 229 39 0,957 0,957 0,957 0,957 0,747 8,048 0,747
7 31312 3 498 169 0,880 0,853 0,880 0,854 4,961 18,680 4,961
8 31552 3 111 13 0,957 0,917 0,957 0,936 0,310 2,594 0,310
9 32537 7 217 54 0,773 0,705 0,773 0,734 2,249 0,906 2,249
10 33315 10 575 351 0,896 0,865 0,896 0,877 40,149 4,819 40,149
11 36895 14 76 11 0,857 0,857 0,857 0,857 0,826 3,260 1,645
12 37364 4 94 24 0,947 0,953 0,947 0,947 0,480 3,327 0,480
13 39582 6 566 352 0,816 0,828 0,816 0,816 24,821 20,419 24,821
14 39716 3 53 24 1,000 1,000 1,000 1,000 0,299 1,468 0,299
15 44077 4 226 18 1,000 1,000 1,000 1,000 0,711 5,468 0,711
The four classification algorithms reveal distinct trade-offs between predictive performance and computational efficiency when applied to high-dimensional gene expression data. For instance, on dataset E-GEOD-20685 (ID 1), SVM attained an accuracy of 95.5% with 107 selected features and a total processing time of 112.137 seconds, while Random Forest achieved a marginally higher accuracy of 97.0% but at a prohibitive total time of 16,306.229 seconds. In contrast, XGBoost matched SVM’s accuracy (95.5%) with a considerably lower training time of 0.395 seconds and an overall time of 40.015 seconds, suggesting a more balanced performance. Gradient Boosting, on the other hand, reached an accuracy of 93.9% using 468 features, with a total time of 17.957 seconds—indicating competitive speed but slightly reduced accuracy compared to the other methods.
Similar trends are observed in other datasets; for example, in E-GEOD-29354 (ID 5), all algorithms achieved perfect accuracy, yet their feature counts and processing times varied substantially. These differences underscore the critical need for an approach that not only maintains high classification performance but also minimizes computational overhead. Although Random Forest often produces high accuracy, its excessive time cost can limit its practicality in real-world applications. Conversely, XGBoost offers a compelling balance between accuracy and efficiency, making it particularly well-suited for gene expression classification tasks. Overall, these results emphasize that algorithm selection should be guided by both predictive metrics and computational feasibility, especially in scenarios involving high-dimensional data.
4.4 Discussion
5 Conclusion and future work
In conclusion, our study introduces BOLIMES, a novel feature selection algorithm that integrates the robustness of Boruta with the interpretability of LIME to tackle the challenges inherent in gene expression classification. By systematically filtering out non-informative genes and ranking the remaining ones based on their local importance, BOLIMES effectively reduces the dimensionality of genomic data while mitigating the risk of overfitting. The iterative evaluation process further refines the feature subset by identifying the optimal number of genes that maximize predictive accuracy, thereby ensuring both high classification performance and enhanced interpretability.
Looking ahead, several promising avenues exist for further advancement of this work. Future research could explore the incorporation of additional interpretability methods alongside LIME to offer a more comprehensive assessment of feature relevance. Moreover, extending BOLIMES to accommodate multi-omics data, such as proteomic and metabolomic profiles, may broaden its applicability in complex biological analyses. Integrating deep learning-based classifiers could also be investigated to further boost performance in intricate gene expression scenarios. Finally, validating the approach on larger, more diverse datasets and real-world clinical samples will be crucial to establish its robustness and generalizability in practical settings.
Acknowledgements: This study is received support from the European Union’s Horizon research and innovation program under the MSCA-SE (Marie Skłodowska-Curie Actions Staff Exchange) grant agreement 101086252; Call: HORIZON-MSCA-2021-SE-01; Project title: STARWARS.

### v2

@@
 Additionally, the high dimensionality increases computational complexity, making conventional classification models inefficient and impractical for real-world applications.
-To address these challenges, feature selection plays a critical role in gene expression classification.
+To address these challenges, feature selection [28, 30] plays a critical role in gene expression classification.
 By identifying and retaining only the most informative genes, feature selection not only improves model generalization but also enhances interpretability, allowing researchers to derive biologically meaningful insights from machine learning predictions [10, 21].
@@
 Unlike conventional methods that rely solely on statistical ranking or classifier-specific selection, BOLIMES integrates the robustness of Boruta [25, 39] with the interpretability of LIME [33], ensuring that only the most relevant and influential genes are retained.
-BOLIMES first applies Boruta, a powerful wrapper method that iteratively refines feature sets by comparing real genes against randomly permuted shadow features, ensuring that no informative gene is prematurely discarded.
-Then, LIME [30, 23] (Local Interpretable Model-agnostic Explanations) is employed to assess the local importance of each selected gene, providing an interpretable ranking of features that capture model-specific contributions.
-Finally, an iterative classification evaluation determines the optimal feature subset by selecting the number of genes that maximize predictive accuracy.
-By integrating Boruta’s exhaustive feature selection with LIME’s interpretability-driven refinement, BOLIMES effectively balances dimensionality reduction with model performance, making it a powerful solution for high-dimensional gene expression analysis.
-The remainder of this paper is structured as follows: we briefly presents a fundamental background in Section 2.
+BOLIMES begins by applying Boruta, safeguarding against the premature elimination of any informative genes.
+Next, LIME (Local Interpretable Model-agnostic Explanations) [33, 24] is employed to assess the local importance of each selected gene, offering an interpretable ranking that highlights the contributions of individual features to the model.
+Finally, an iterative classification evaluation identifies the optimal subset of features by selecting the number of genes that maximize predictive accuracy.
+We pay attention at Boruta and LIME for the following reasons: (1) Due to the large number of dimensions, LIME would face scalability issues when it comes to interpretation.
+(2) Boruta is viewed as a global feature selection method, whereas LIME focuses on local feature selection.
+Therefore, we will first perform global filtering using Boruta, followed by local refinement with LIME.
+(3) Our primary objective is to determine the optimal number of features for selection.
+As a result, we have proposed the Bolimex algorithm to select the most appropriate set of features, ensuring the best possible accuracy.
+The remainder of this paper [^1] This paper will be condensed to meet the page limit requirements of the conference upon acceptance.
+We aim to retain all the core ideas and details from this version to ensure the integrity of our research.
+is structured as follows: we briefly presents a fundamental background in Section 2.
 Next, Section 3 describes our proposal algorithm.
@@
 2 Background
-In this section, we present the foundation of our approach with integrating Boruta and Lime. We also provide the ML algorithm that will be used for the gene expression classification task.
+In this section, we present the foundation of our approach with integrating Boruta and Lime. We also provide the ML models that will be used for the gene expression classification task.
 2.1 Gene Expression Classification with ML models
@@
 Once trained, $f$ is used to classify new, unseen inputs.
-While numerous feature selection techniques exist, our study concentrates on two: Boruta and LIME.
-We choose Boruta for its robust ability to identify all truly relevant features within high-dimensional datasets, ensuring that no significant predictor is overlooked.
-Complementing this, LIME is employed for its exceptional capacity to provide local, interpretable explanations of model predictions, which is crucial when assessing feature importance through numerous distance calculations.
-Together, these methods offer a powerful and efficient approach to refining our gene expression classification models.
-Now, we introduce Boruta and Lime in the next session.
+Feature selection is crucial before classification begins. Our study focuses on two techniques: Boruta and LIME. We now introduce Boruta and LIME in the following sections.
 2.2 Leveraging Boruta for Robust Feature Extraction
@@
 This rigorous selection process is particularly valuable in high-dimensional applications, such as gene expression classification, where capturing meaningful signals is crucial.
-Rather than directly improving predictive accuracy, Boruta refines the feature set, which can indirectly enhance model interpretability and performance.
 For clarity, we formally define Boruta as follows:
@@
   4. 4.
-    Iteration: Remove irrelevant features and repeat until all features are decisively classified.
+    Iteration: Remove irrelevant and tentative features and repeat until all features are decisively classified.
 The final selected subset $X^{*}\subseteq X$ comprises all features deemed relevant.
+After applying the Boruta algorithm, we retain only the relevant features (confirmed) and excluded the tentative and irrelevant features (rejected).
+To further enhance the selection of features in $X^{*}$, we employed the AI explanation technique outlined in the following section.
 2.3 XAI for Feature Selection
 Explainable AI (XAI) [13, 38] represents a forefront of AI research, aiming to elucidate the decision-making processes of complex models.
-In the context of gene expression classification, where feature selection is pivotal to model performance and interpretability, our study leverages LIME—Local Interpretable Model-Agnostic Explanations—to demystify and extract critical features.
+In the context of gene expression classification, where feature selection is pivotal to model performance and interpretability, our study leverages LIME—Local Interpretable Model-Agnostic Explanations—to demystify and select critical features.
 LIME approximates the behavior of a sophisticated, black-box model with a simpler, locally interpretable surrogate, thereby pinpointing the most influential predictors in the vicinity of a given instance.
@@
 The absolute coefficients $|w_{j}|$ indicate the local importance of each feature, enabling a further refined selection $X_{\text{selected}}\subseteq X^{*}$ for classification.
-3 BOLIMEX algorithm
+To clarify, our choice of LIME for feature selection arises from the critical question of determining the optimal number of features for the model.
+In this context, assessing the local importance of each vector proves to be the most effective strategy, leading us to introduce the BOLIMES algorithm.
+The following section will provide a comprehensive explanation of the BOLIMES algorithm and its application.
+3 BOLIMES algorithm
 Our methodology begins with applying the Boruta algorithm to sift through the high-dimensional gene expression dataset, effectively filtering out irrelevant features and isolating those that are truly significant.
 This initial reduction is critical because LIME, our subsequent interpretability tool, involves generating numerous perturbed samples and calculating distances—a computationally intensive process, especially in large feature spaces.
-By narrowing the focus to a refined subset via Boruta, we significantly reduce the computational burden and enhance the precision of LIME’s local explanations.
-This sequential approach streamlines the overall feature selection process and fortifies the reliability and clarity of our model’s interpretability, ultimately leading to improved classification performance.
+By narrowing the feature set, we improve LIME’s precision and enhance the model’s interpretability, ultimately boosting classification performance.
 However, the key question remains: how many features should be selected for optimal classification?
@@
 8
-9      for $k=1$ to $|X^{*}|$ do
+9      for $k=10$ to $|X^{*}|$ do
              $S_{k}\leftarrow\{x^{*}_{(i)}\}_{i=1}^{k}$
@@
 Algorithm 1 Optimal Feature Selection for Classification using Boruta and LIME
-The algorithm presents an optimal feature selection framework by integrating Boruta and LIME to refine feature subsets for classification.
+Algorithm 1 presents an optimal feature selection framework by integrating Boruta and LIME to refine feature subsets for classification.
 Given a dataset $D=(X,y)$, it first applies Boruta to extract a subset $X^{*}$ of relevant features.
@@
 $S_{k}\leftarrow\{x^{*}_{(i)}\}_{i=1}^{k},\quad f_{k}\leftarrow\text{TrainClassifier}(D_{S_{k}}).$
+Note that we begin with $k=10$ and increase it until $|X^{*}|$, as using a smaller number of vectors would be inefficient.
 The best-performing classifier defines $k^{*}$, yielding the final subset $X_{\text{opt}}$ and trained model $f_{\text{opt}}$.
@@
 Furthermore, we also provide the results of Boruta.
-Our source code has been made publicly accessible on GitHub [^1] https://github.com/Vannguyen0312/Manholes-App/.
+Our source code has been made publicly accessible on GitHub [^2] https://github.com/Pbchung75/BOLIMES/.
 4.1 Dataset and Configurations
-The gene expression datasets summarized in Table 1 exemplify the inherent challenges of high-dimensional biomedical data.
+We carry out a series of experiments using 14 gene expression datasets, all meticulously sourced from the reputable ArrayExpress repository [4].
+The gene expression datasets [10] summarized in Table 1 exemplify the inherent challenges of high-dimensional biomedical data.
 With sample sizes ranging from $53$ to $575$ and feature counts spanning from approximately $11,950$ to over $54,600$, these datasets present a significant imbalance between the number of available samples and the vast dimensionality of gene expression profiles.
-Additionally, the variability in the number of classes—from as few as 3 to as many as 14—further complicates the classification task by introducing diverse and complex biological signals.
+Additionally, the variability in the number of classes—from as few as $3$ to as many as $10$—further complicates the classification task by introducing diverse and complex biological signals.
 This high dimensionality coupled with limited sample sizes accentuates the risk of overfitting and underscores the critical need for effective feature selection.
-Robust feature selection methods are essential to isolate the most informative genes, thereby enhancing model interpretability and predictive accuracy in gene expression classification.
 In this study, our feature selection strategy is specifically designed to address these challenges, ensuring that only the most relevant features are retained for subsequent classification tasks.
@@
 10 E-GEOD-33315 575 22283 10
-11 E-GEOD-36895 76 54675 14
 11 E-GEOD-37364 94 54675 4
@@
 14 E-GEOD-44077 226 33252 4
-To evaluate our approach and the training model, we implement the AI libraries (i.e., tensorflow (version 3.10.0), Keras (version 3.0)) and run the experiments on the computer with the following configuration: AMD Ryzen 7 6800HS, Ceator Edition 3.20 GHZ, 16 GB RAM, Windows 10 OS.
+To evaluate our approach and the training model, we implement the AI libraries (i.e., Pandas (version 2.2.2), Scikit-learn (version 1.6.1)) and run the experiments on the computer with the following configuration: Intel Core i5-12400, 2.50 GHz, 24 GB RAM, Windows 11 Pro OS.
 4.2 Boruta’s Feature Selection
-In this study, we adopt Boruta as our primary feature selection method to substantially reduce the dimensionality of our gene expression datasets before applying LIME for model interpretation.
+In this study, we adopt Boruta as our primary feature selection method to substantially reduce the dimensionality of our gene expression datasets for the global problem.
 The rationale behind this choice is to prevent an explosion in computational complexity and potential loss of interpretability when LIME is applied to an excessively high-dimensional feature space.
-By using Boruta, we are able to effectively eliminate irrelevant features while retaining those that are truly informative for classification.
-As evidenced in Table 2, datasets such as E-GEOD-20685 were reduced from 54,627 dimensions to only 545 confirmed features, thereby rendering the subsequent LIME analysis both feasible and efficient.
-To achieve this, Boruta was executed with the following parameters via the BorutaPy library: a random forest classifier (‘rf‘) with 300 estimators, a maximum of 200 iterations, an alpha value of ’0.01’, a percentile threshold of 95, two-step feature selection enabled, and a fixed random state of 42, with verbose output enabled.
-These parameter settings were meticulously chosen to ensure a rigorous and robust selection process, ultimately facilitating a more interpretable and high-performing classification model.
+By using Boruta, we are able to effectively eliminate irrelevant features (tentative and rejected in Table 2) while retaining those that are truly informative for classification.
+As evidenced in Table 2, datasets such as E-GEOD-20685 are reduced from 54,627 dimensions to only 545 confirmed features, thereby rendering the subsequent LIME analysis both feasible and efficient.
+To achieve this, Boruta is executed with the following parameters via the BorutaPy library: a random forest classifier (‘rf‘) with 300 estimators, a maximum of 200 iterations, an alpha value of $0.01$, a strict percentile threshold of 100, two-step feature selection enabled, and a fixed random state of 42, with verbose output enabled.
+These parameter settings are meticulously selected to ensure a rigorous and robust selection process, ultimately facilitating a more interpretable and high-performing classification model.
 Table 2: Feature Selection Results
@@
 10 33315 483 108 21692 1176,811348
-11 36895 39 14 54622 122,578902
 11 37364 59 43 54573 129,0980539
@@
 4.3 Classification results
-In our study, we conducted extensive experiments with four advanced algorithms to determine the method yielding the highest accuracy across various gene expression datasets.
-We configured an SVM using SVC with a radial basis function kernel, setting C to 100,000, gamma to ”scale”, and class weights to ”balanced”, ensuring reproducibility with a random state of 42.
-In parallel, we employed a Random Forest classifier with 200 estimators, a maximum depth of 10, and balanced class weights.
-Additionally, we implemented both an XGBClassifier and a GradientBoostingClassifier, each configured with 50 estimators, a maximum depth of 10, and a learning rate of 0.01, with the random state fixed at 42.
-Our comparative evaluation revealed that while the SVM excelled on datasets with fewer samples, the ensemble methods—particularly Random Forest and XGBoost—demonstrated more robust performance on high-dimensional data.
-This comprehensive analysis highlights the critical role of algorithm selection and fine-tuning in achieving optimal classification performance in gene expression studies.
+In this study, we conduct experiments with four ML algorithms (using SVM, Random Forest (RF), XGBoost (XGB), and Gradient Boosting (GB)) to determine the method yielding the highest accuracy across various gene expression datasets (see Table 3, 4, 5, and 6).
+We configure an SVM using SVC with a radial basis function kernel, setting C to $100,000$, gamma to $0.001$, ensuring reproducibility with a random state of $42$.
+In parallel, we employe a Random Forest classifier with $200$ estimators, a maximum depth of $10$.
+Additionally, we implemented both an XGBClassifier and a GradientBoostingClassifier, each configured with $50$ estimators, a maximum depth of $10$, and a learning rate of $0.01$.
+Our comparative evaluation reveales that while the SVM excelled on datasets with fewer samples, the ensemble methods—particularly Random Forest and XGBoost—demonstrated more robust performance on high-dimensional data.
+In Table 3, 4, 5, and 6, we conduct SVM, RF, and Decision Tree (DT) methodology to identify the importance of features and the number of selected features ($k$ selected dimensions) following the methodology outlined in [10] i.e., Top $k$ selected feature of E-GEOD-20685 (ID 1) is $85$, E-GEOD-20711 (ID 2) is $43$, E-GEOD-21050 (ID 3) is $275$, and others.
+Overall, our proposed algorithm demonstrated superior performance, achieving higher accuracy with respect to the F1-score.
+In certain instances, it reached $100\%$, while other solutions (SVM, RF, DT) only achieved $80-90\%$.
+The feature selection algorithms, such as SVM, RF, and DT, with a smaller number of features, appeared to underperform in covering the data (e.g., IDs 1, 2, 4).
+Our approach, however, seemed to be more effective in determining the optimal number of features needed.
 Table 3: SVM Classification Results with Top $k$ Features
-ID Dataset Class Method Samples Top K Classification Results Time (s)
-Features Acc Prec Rec F1 Score Training Preprocess Total
-ours 327 107 0,955 0,956 0,955 0,955 1,151 12,951 112,137
-1 20685 6 327 107 0,955 0,956 0,955 0,955 1,151 12,951 112,137
-327 107 0,955 0,956 0,955
-2 20711 5 90 67 0,944 0,952 0,944 0,944 0,298 3,745 17,754
-3 21050 4 310 70 0,677 0,710 0,677 0,678 1,494 10,400 295,982
-4 21122 7 158 170 0,938 0,915 0,938 0,922 0,548 2,010 82,092
-5 29354 3 53 10 1,000 1,000 1,000 1,000 0,165 0,823 0,165
-6 30784 3 229 14 1,000 1,000 1,000 1,000 0,730 8,197 3,476
-7 31312 3 498 43 0,900 0,900 0,900 0,900 0,063 19,236 2,144
-8 31552 3 111 15 0,957 0,917 0,957 0,936 0,299 2,642 1,816
-9 32537 7 217 40 0,796 0,730 0,796 0,747 0,735 0,918 21,201
-10 33315 10 575 273 0,878 0,860 0,878 0,867 0,138 4,844 33,250
-11 36895 14 76 15 0,929 0,893 0,929 0,905 0,228 3,199 1,304
-12 37364 4 94 17 0,947 0,965 0,947 0,947 0,259 3,393 2,083
-13 39582 6 566 251 0,877 0,889 0,877 0,880 0,130 20,512 25,058
-14 39716 3 53 12 1,000 1,000 1,000 1,000 0,157 1,642 0,454
-15 44077 4 226 18 1,000 1,000 1,000 1,000 0,578 5,513 5,220
-The classification results obtained using SVM, Random Forest, XGBoost, and Gradient Boosting reveal distinct performance and computational trade-offs.
-SVM generally delivers high accuracy and balanced precision, recall, and F1 scores across most datasets, with relatively low training times in some cases; however, its preprocessing time can be substantial when handling a large number of features.
-In contrast, Random Forest exhibits competitive accuracy, often outperforming SVM on certain datasets, yet it tends to incur significantly higher training times, particularly when the number of top features increases—as observed in dataset 1 where training time was notably high.
-XGBoost demonstrates robust performance with competitive accuracy and precision; its moderate training and preprocessing times suggest that it strikes a balance between model complexity and computational efficiency, though its performance can be dataset-dependent (e.g., lower accuracy on dataset 13).
-Meanwhile, Gradient Boosting achieves near-perfect accuracy on several datasets, indicating high robustness in classification, yet its training and total processing times vary considerably across datasets.
+ID Dataset Class Method Samples Top k Classification Results Training
+Features Acc Prec Rec F1 Score Time (s)
+1 20685 6 ours 327 107 0.955 0.956 0.955 0.955 112.137
+SVM 85 0.859 0.872 0.859 0.859 0.135
+RF 85 0.890 0.902 0.890 0.889 0.147
+DT 85 0.872 0.884 0.872 0.868 0.192
+2 20711 5 ours 90 67 0.944 0.952 0.944 0.943 17.754
+SVM 43 0.811 1.000 0.811 0.811 0.818
+RF 43 0.778 1.000 0.778 0.778 0.806
+DT 43 0.778 1.000 0.778 0.778 0.788
+3 21050 4 ours 310 70 0.677 0.709 0.677 0.678 295.982
+SVM 275 0.684 0.690 0.684 0.676 0.218
+RF 275 0.681 0.681 0.681 0.666 0.227
+DT 275 0.652 0.674 0.652 0.622 0.250
+4 21122 7 ours 158 170 0.938 0.915 0.938 0.922 82.092
+SVM 78 0.886 1.000 0.886 0.886 1.489
+RF 78 0.880 1.000 0.880 0.880 1.483
+DT 78 0.842 1.000 0.842 0.842 1.744
+5 29354 3 ours 53 10 1.000 1.000 1.000 1.000 0.165
+SVM 35 0.868 1.000 0.868 0.868 0.415
+RF 35 0.830 1.000 0.830 0.830 0.422
+DT 35 0.660 1.000 0.660 0.660 0.397
+6 30784 3 ours 229 14 1.000 1.000 1.000 1.000 3.476
+SVM 42 0.904 1.000 0.904 0.904 2.277
+RF 42 0.908 1.000 0.908 0.908 2.075
+DT 42 0.908 1.000 0.908 0.908 2.175
+7 31312 3 ours 498 43 0.900 0.900 0.900 0.900 2.144
+SVM 195 0.878 0.872 0.878 0.859 0.236
+RF 195 0.878 0.864 0.878 0.857 0.225
+DT 195 0.809 0.796 0.809 0.775 0.345
+8 31552 3 ours 111 15 0.957 0.917 0.957 0.936 1.816
+SVM 44 0.820 1.000 0.820 0.820 0.931
+RF 44 0.883 1.000 0.883 0.883 0.870
+DT 44 0.847 1.000 0.847 0.847 0.890
+9 32537 7 ours 217 40 0.795 0.730 0.795 0.747 21.201
+SVM 171 0.760 1.000 0.760 0.760 3.054
+RF 171 0.779 1.000 0.779 0.779 2.910
+DT 171 0.756 1.000 0.756 0.756 3.450
+10 33315 10 ours 575 273 0.878 0.860 0.878 0.867 33.250
+SVM 2321 0.873 0.893 0.873 0.855 4.174
+RF 2321 0.870 0.893 0.870 0.851 4.667
+DT 2321 0.781 0.845 0.781 0.752 5.534
+11 37364 4 ours 94 17 0.947 0.965 0.947 0.947 2.083
+SVM 140 0.809 1.000 0.809 0.809 0.793
+RF 140 0.755 1.000 0.755 0.755 0.750
+DT 140 0.713 1.000 0.713 0.713 0.798
+12 39582 6 ours 566 251 0.877 0.889 0.877 0.880 25.058
+SVM 441 0.859 0.873 0.859 0.859 0.466
+RF 441 0.862 0.874 0.862 0.861 0.480
+DT 441 0.790 0.812 0.790 0.786 0.764
+13 39716 3 ours 53 12 1.000 1.000 1.000 1.000 0.454
+SVM 118 0.887 1.000 0.887 0.887 0.408
+RF 118 0.925 1.000 0.925 0.925 0.408
+DT 118 0.774 1.000 0.774 0.774 0.401
+14 44077 4 ours 226 18 1.000 1.000 1.000 1.000 5.220
+SVM 23 0.991 1.000 0.991 0.991 1.803
+RF 23 0.996 1.000 0.996 0.996 1.790
+DT 23 0.987 1.000 0.987 0.987 2.005
 Table 4: Random Forest Classification Results with Top $k$ Features
-ID Dataset Class Samples Top N Classification Results Time (s)
-Features Acc Prec Rec F1 Score Training Preprocess Total
-1 20685 6 327 487 0,970 0,971 0,970 0,970 40,762 13,055 16306,229
-2 20711 5 90 10 0,833 0,841 0,833 0,830 5,396 3,703 5,396
-3 21050 4 310 72 0,726 0,727 0,726 0,711 25,093 10,252 1507,624
-4 21122 7 158 104 0,938 0,922 0,938 0,925 11,577 1,971 1039,209
-5 29354 3 53 10 1,000 1,000 1,000 1,000 3,325 1,059 3,325
-6 30784 3 229 12 1,000 1,000 1,000 1,000 14,531 8,278 43,917
-7 31312 3 498 161 0,900 0,821 0,900 0,858 1,594 19,739 221,322
-8 31552 3 111 12 1,000 1,000 1,000 1,000 6,667 2,608 20,318
-9 32537 7 217 87 0,773 0,680 0,773 0,707 17,146 0,935 1238,228
-10 33315 10 575 444 0,887 0,854 0,887 0,868 3,215 4,947 985,866
-11 36895 14 76 20 0,929 0,929 0,929 0,929 4,347 3,242 46,355
-12 37364 4 94 10 0,947 0,953 0,947 0,947 5,579 3,380 5,579
-13 39582 6 566 189 0,877 0,890 0,877 0,876 1,908 20,970 304,137
-14 39716 3 53 15 1,000 1,000 1,000 1,000 3,172 1,528 19,338
-15 44077 4 226 52 1,000 1,000 1,000 1,000 15,408 6,125 642,643
+ID Dataset Class Method Samples Top k Classification Results Training
+Features Acc Prec Rec F1 Score Time (s)
+1 20685 6 Ours 327 48 0.970 0.971 0.970 0.970 16306.229
+SVM 85 0.838 0.852 0.838 0.831 500.484
+RF 85 0.874 0.890 0.874 0.872 4.706
+DT 85 0.887 0.904 0.887 0.885 5.175
+2 20711 5 Ours 90 10 0.833 0.841 0.833 0.830 5.396
+SVM 43 0.767 1.000 0.767 0.767 25.742
+RF 43 0.833 1.000 0.833 0.833 2593.662
+DT 43 0.833 1.000 0.833 0.833 2588.751
+3 21050 4 Ours 310 72 0.726 0.727 0.726 0.711 1507.624
+SVM 275 0.732 0.776 0.732 0.696 6.371
+RF 275 0.735 0.773 0.735 0.700 6.617
+DT 275 0.739 0.793 0.739 0.701 6.441
+4 21122 7 Ours 158 104 0.938 0.922 0.938 0.925 1039.209
+SVM 78 0.880 1.000 0.880 0.880 50.215
+RF 78 0.886 1.000 0.886 0.886 50.446
+DT 78 0.873 1.000 0.873 0.873 53.083
+5 29354 3 Ours 53 10 1.000 1.000 1.000 1.000 3.325
+SVM 35 0.887 1.000 0.887 0.887 13.873
+RF 35 0.811 1.000 0.811 0.811 12.891
+DT 35 0.774 1.000 0.774 0.774 12.654
+6 30784 3 Ours 229 12 1.000 1.000 1.000 1.000 43.917
+SVM 42 0.913 1.000 0.913 0.913 80.499
+RF 42 0.895 1.000 0.895 0.895 75.487
+DT 42 0.908 1.000 0.908 0.908 76.522
+7 31312 3 Ours 498 161 0.900 0.821 0.900 0.858 221.322
+SVM 195 0.871 0.885 0.871 0.828 8.221
+RF 195 0.872 0.885 0.872 0.828 7.968
+DT 195 0.841 0.859 0.841 0.800 8.715
+8 31552 3 Ours 111 12 1.000 1.000 1.000 1.000 20.318
+SVM 44 0.865 1.000 0.865 0.865 31.843
+RF 44 0.901 1.000 0.901 0.901 29.352
+DT 44 0.865 1.000 0.865 0.865 30.185
+9 32537 7 Ours 217 87 0.773 0.679 0.773 0.707 1238.228
+SVM 171 0.779 1.000 0.779 0.779 113.423
+RF 171 0.793 1.000 0.793 0.793 104.033
+DT 171 0.765 1.000 0.765 0.765 110.804
+10 33315 10 Ours 575 444 0.887 0.854 0.887 0.868 985.966
+SVM 2321 0.835 0.866 0.835 0.808 31.634
+RF 2321 0.833 0.868 0.833 0.800 31.258
+DT 2321 0.797 0.840 0.797 0.756 31.688
+11 37364 4 Ours 94 10 0.947 0.953 0.947 0.947 5.579
+SVM 140 0.798 1.000 0.798 0.798 26.410
+RF 140 0.809 1.000 0.809 0.809 25.745
+DT 140 0.734 1.000 0.734 0.734 27.083
+12 39582 6 Ours 566 189 0.877 0.890 0.877 0.876 304.137
+SVM 441 0.820 0.836 0.820 0.819 13.926
+RF 441 0.832 0.847 0.832 0.831 13.629
+DT 441 0.807 0.831 0.807 0.805 14.261
+13 39716 3 Ours 53 15 1.000 1.000 1.000 1.000 19.338
+SVM 118 0.925 1.000 0.925 0.925 13.133
+RF 118 0.925 1.000 0.925 0.925 13.049
+DT 118 0.830 1.000 0.830 0.830 13.391
+14 44077 4 Ours 226 52 1.000 1.000 1.000 1.000 642.643
+SVM 23 0.991 1.000 0.991 0.991 61.789
+RF 23 0.996 1.000 0.996 0.996 60.206
+DT 23 0.996 1.000 0.996 0.996 63.950
 Table 5: XGB Classification Results with Top $k$ Features
-IID Dataset Class Samples Top N Classification Results Time (s)
-Features Acc Prec Rec F1 Score Training Preprocess Total
-1 20685 6 327 114 0,955 0,962 0,955 0,955 0,395 14,439 40,015
-2 20711 5 90 10 0,833 0,841 0,833 0,830 0,328 3,800 0,328
-3 21050 4 310 60 0,790 0,791 0,790 0,784 0,286 11,975 16,186
-4 21122 7 158 159 0,938 0,921 0,938 0,925 0,293 4,028 44,087
-5 29354 3 53 10 1,000 1,000 1,000 1,000 0,271 2,541 0,271
-6 30784 3 229 50 0,978 0,979 0,978 0,976 0,127 10,126 6,344
-7 31312 3 498 10 0,860 0,848 0,860 0,843 0,388 21,554 0,388
-8 31552 3 111 10 1,000 1,000 1,000 1,000 0,557 4,239 0,557
-9 32537 7 217 95 0,796 0,693 0,796 0,736 0,299 2,478 26,388
-10 33315 10 575 425 0,870 0,842 0,870 0,855 1,341 6,495 389,225
-11 36895 14 76 29 0,929 0,929 0,929 0,929 0,250 3,364 6,509
-12 37364 4 94 21 1,000 1,000 1,000 1,000 0,142 4,804 1,862
-13 39582 6 566 556 0,825 0,833 0,825 0,824 1,564 22,414 554,762
-14 39716 3 53 44 1,000 1,000 1,000 1,000 0,101 3,074 3,746
-15 44077 4 226 57 0,978 0,980 0,978 0,979 0,139 6,925 8,332
-Table 6: Gradient Boosting Classification Results with Top N Features
-IID Dataset Class Samples Top N Classification Results Time (s)
-Features Acc Prec Rec F1 Score Training Preprocess Total
-1 20685 6 327 468 0,939 0,947 0,939 0,936 17,957 12,795 17,957
-2 20711 5 90 38 0,889 0,921 0,889 0,892 0,664 3,832 0,664
-3 21050 4 310 65 0,807 0,802 0,807 0,804 1,844 9,923 1,844
-4 21122 7 158 257 0,875 0,872 0,875 0,871 5,544 1,890 5,544
-5 29354 3 53 12 1,000 1,000 1,000 1,000 0,387 0,822 0,387
-6 30784 3 229 39 0,957 0,957 0,957 0,957 0,747 8,048 0,747
-7 31312 3 498 169 0,880 0,853 0,880 0,854 4,961 18,680 4,961
-8 31552 3 111 13 0,957 0,917 0,957 0,936 0,310 2,594 0,310
-9 32537 7 217 54 0,773 0,705 0,773 0,734 2,249 0,906 2,249
-10 33315 10 575 351 0,896 0,865 0,896 0,877 40,149 4,819 40,149
-11 36895 14 76 11 0,857 0,857 0,857 0,857 0,826 3,260 1,645
-12 37364 4 94 24 0,947 0,953 0,947 0,947 0,480 3,327 0,480
-13 39582 6 566 352 0,816 0,828 0,816 0,816 24,821 20,419 24,821
-14 39716 3 53 24 1,000 1,000 1,000 1,000 0,299 1,468 0,299
-15 44077 4 226 18 1,000 1,000 1,000 1,000 0,711 5,468 0,711
+ID Dataset Class Method Samples Top k Classification Results Training
+Features Acc Prec Rec F1 Score Time (s)
+1 20685 6 Ours 327 114 0.955 0.961 0.955 0.955 40.015
+SVM 85 0.804 0.839 0.804 0.793 10.932
+RF 85 0.829 0.854 0.829 0.822 11.592
+DT 85 0.851 0.874 0.851 0.843 11.531
+2 20711 5 Ours 90 10 0.833 0.841 0.833 0.830 0.328
+SVM 43 0.744 1.000 0.744 0.744 62.459
+RF 43 0.811 1.000 0.811 0.811 63.819
+DT 43 0.811 1.000 0.811 0.811 63.947
+3 21050 4 Ours 310 60 0.790 0.791 0.790 0.784 16.183
+SVM 275 0.732 0.769 0.732 0.701 9.478
+RF 275 0.748 0.782 0.748 0.724 9.977
+DT 275 0.752 0.796 0.752 0.720 9.622
+4 21122 7 Ours 158 159 0.938 0.921 0.938 0.924 44.087
+SVM 78 0.848 1.000 0.848 0.848 162.044
+RF 78 0.854 1.000 0.854 0.854 164.271
+DT 78 0.867 1.000 0.867 0.867 166.522
+5 29354 3 Ours 53 10 1.000 1.000 1.000 1.000 0.271
+SVM 35 0.906 1.000 0.906 0.906 23.871
+RF 35 0.774 1.000 0.774 0.774 23.896
+DT 35 0.830 1.000 0.830 0.830 23.359
+6 30784 3 Ours 229 50 0.978 0.979 0.978 0.976 6.344
+SVM 42 0.921 1.000 0.921 0.921 121.816
+RF 42 0.913 1.000 0.913 0.913 119.124
+DT 42 0.921 1.000 0.921 0.921 123.405
+7 31312 3 Ours 498 10 0.860 0.848 0.860 0.843 0.388
+SVM 195 0.851 0.863 0.851 0.810 7.677
+RF 195 0.833 0.828 0.833 0.797 7.432
+DT 195 0.831 0.850 0.831 0.790 7.762
+8 31552 3 ours 111 10 1.000 1.000 1.000 1.000 0.557
+SVM 44 0.856 1.000 0.856 0.856 53.782
+RF 44 0.892 1.000 0.892 0.892 52.092
+DT 44 0.892 1.000 0.892 0.892 53.220
+9 32537 7 ours 217 95 0.795 0.693 0.795 0.736 26.388
+SVM 171 0.783 1.000 0.783 0.783 282.386
+RF 171 0.802 1.000 0.802 0.802 283.478
+DT 171 0.770 1.000 0.770 0.770 283.282
+10 33315 10 ours 575 425 0.870 0.842 0.870 0.854 389.225
+SVM 2321 0.861 0.883 0.861 0.838 94.119
+RF 2321 0.859 0.879 0.859 0.839 94.236
+DT 2321 0.852 0.874 0.852 0.827 96.417
+11 37364 4 ours 94 21 1.000 1.000 1.000 1.000 1.862
+SVM 140 0.777 1.000 0.777 0.777 59.093
+RF 140 0.798 1.000 0.798 0.798 59.041
+DT 140 0.745 1.000 0.745 0.745 60.416
+12 39582 6 ours 566 556 0.825 0.833 0.825 0.824 554.762
+SVM 441 0.797 0.810 0.797 0.793 21.655
+RF 441 0.804 0.816 0.804 0.799 22.184
+DT 441 0.788 0.791 0.788 0.784 22.569
+13 39716 3 ours 53 44 1.000 1.000 1.000 1.000 3.746
+SVM 118 0.897 1.000 0.887 0.887 23.631
+RF 118 0.925 1.000 0.925 0.925 22.742
+DT 118 0.943 1.000 0.943 0.943 24.227
+14 44077 4 ours 226 57 0.978 0.980 0.978 0.978 8.332
+SVM 23 0.978 1.000 0.978 0.978 132.205
+RF 23 0.987 1.000 0.987 0.987 133.960
+DT 23 0.991 1.000 0.991 0.991 133.546
+Table 6: Gradient Boosting Classification Results with Top $k$ Features
+ID Dataset Class Method Samples Top k Classification Results Training
+Features Acc Prec Rec F1 Score Time (s)
+1 20685 6 ours 327 468 0.939 0.947 0.939 0.936 17.957
+SVM 85 0.795 0.820 0.795 0.788 53.840
+RF 85 0.856 0.876 0.856 0.855 53.966
+DT 85 0.844 0.865 0.844 0.834 53.381
+2 20711 5 ours 90 38 0.889 0.921 0.889 0.892 0.664
+SVM 43 0.733 1.000 0.733 0.733 91.394
+RF 43 0.711 1.000 0.711 0.711 90.162
+DT 43 0.711 1.000 0.711 0.711 89.505
+3 21050 4 ours 310 65 0.806 0.802 0.806 0.804 1.844
+SVM 275 0.703 0.733 0.703 0.685 72.745
+RF 275 0.735 0.745 0.735 0.720 78.348
+DT 275 0.758 0.792 0.758 0.742 76.235
+4 21122 7 ours 158 257 0.875 0.872 0.875 0.871 5.544
+SVM 78 0.804 1.000 0.804 0.804 511.805
+RF 78 0.835 1.000 0.835 0.835 507.856
+DT 78 0.861 1.000 0.861 0.861 513.399
+5 29354 3 ours 53 12 1.000 1.000 1.000 1.000 0.387
+SVM 35 0.849 1.000 0.849 0.849 22.807
+RF 35 0.717 1.000 0.717 0.717 20.737
+DT 35 0.811 1.000 0.811 0.811 20.797
+6 30784 3 ours 229 39 0.957 0.957 0.957 0.957 0.747
+SVM 42 0.913 1.000 0.913 0.913 272.929
+RF 42 0.900 1.000 0.900 0.900 269.238
+DT 42 0.943 1.000 0.943 0.943 273.324
+7 31312 3 ours 498 169 0.880 0.853 0.880 0.854 4.961
+SVM 195 0.853 0.840 0.853 0.830 89.338
+RF 195 0.853 0.828 0.853 0.824 90.304
+DT 195 0.825 0.805 0.825 0.795 90.659
+8 31552 3 ours 111 13 0.957 0.917 0.957 0.936 0.310
+SVM 44 0.802 1.000 0.802 0.802 77.917
+RF 44 0.883 1.000 0.883 0.883 74.657
+DT 44 0.892 1.000 0.892 0.892 70.156
+9 32537 7 ours 217 54 0.773 0.705 0.773 0.734 2.249
+SVM 171 0.747 1.000 0.747 0.747 1942.324
+RF 171 0.742 1.000 0.742 0.742 1928.460
+DT 171 0.747 1.000 0.747 0.747 1943.035
+10 33315 10 ours 575 351 0.896 0.865 0.896 0.877 40.149
+SVM 2321 0.833 0.847 0.833 0.822 4118.864
+RF 2321 0.856 0.873 0.856 0.839 4055.604
+DT 2321 0.840 0.852 0.840 0.824 3923.718
+11 37364 4 ours 94 24 0.947 0.953 0.947 0.947 0.480
+SVM 140 0.787 1.000 0.787 0.787 173.044
+RF 140 0.745 1.000 0.745 0.745 172.476
+DT 140 0.745 1.000 0.745 0.745 172.407
+12 39582 6 ours 566 352 0.816 0.828 0.816 0.816 24.821
+SVM 441 0.802 0.819 0.802 0.799 465.967
+RF 441 0.813 0.819 0.813 0.806 465.414
+DT 441 0.776 0.794 0.776 0.768 470.792
+13 39716 3 ours 53 24 1.000 1.000 1.000 1.000 0.299
+SVM 118 0.906 1.000 0.906 0.906 38.331
+RF 118 0.868 1.000 0.868 0.868 34.458
+DT 118 0.811 1.000 0.811 0.811 38.443
+14 44077 4 ours 226 18 1.000 1.000 1.000 1.000 0.711
+SVM 23 0.973 1.000 0.973 0.973 220.643
+RF 23 0.982 1.000 0.982 0.982 210.123
+DT 23 0.987 1.000 0.987 0.987 224.643
 The four classification algorithms reveal distinct trade-offs between predictive performance and computational efficiency when applied to high-dimensional gene expression data.
@@
 Conversely, XGBoost offers a compelling balance between accuracy and efficiency, making it particularly well-suited for gene expression classification tasks.
-Overall, these results emphasize that algorithm selection should be guided by both predictive metrics and computational feasibility, especially in scenarios involving high-dimensional data.
-4.4 Discussion
+In general, the results of our proposed method outperformed feature selection techniques in terms of accuracy.
+Moreover, the computational time for the training process was significantly reduced, as our approach identified fewer but more crucial and promising features compared to previous studies.
+This reduction in the number of features also led to a faster training time.
+In some cases, although more features were selected, the accuracy achieved was considerably higher than that of other solutions
 5 Conclusion and future work
-In conclusion, our study introduces BOLIMES, a novel feature selection algorithm that integrates the robustness of Boruta with the interpretability of LIME to tackle the challenges inherent in gene expression classification.
-By systematically filtering out non-informative genes and ranking the remaining ones based on their local importance, BOLIMES effectively reduces the dimensionality of genomic data while mitigating the risk of overfitting.
-The iterative evaluation process further refines the feature subset by identifying the optimal number of genes that maximize predictive accuracy, thereby ensuring both high classification performance and enhanced interpretability.
-Looking ahead, several promising avenues exist for further advancement of this work.
-Future research could explore the incorporation of additional interpretability methods alongside LIME to offer a more comprehensive assessment of feature relevance.
-Moreover, extending BOLIMES to accommodate multi-omics data, such as proteomic and metabolomic profiles, may broaden its applicability in complex biological analyses.
-Integrating deep learning-based classifiers could also be investigated to further boost performance in intricate gene expression scenarios.
-Finally, validating the approach on larger, more diverse datasets and real-world clinical samples will be crucial to establish its robustness and generalizability in practical settings.
-Acknowledgements: This study is received support from the European Union’s Horizon research and innovation program under the MSCA-SE (Marie Skłodowska-Curie Actions Staff Exchange) grant agreement 101086252; Call: HORIZON-MSCA-2021-SE-01; Project title: STARWARS.
+This study presents BOLIMES, a novel feature selection algorithm that combines the robustness of Boruta with the interpretability of LIME to address the challenges in gene expression classification.
+The proposed method has demonstrated superior performance, achieving higher accuracy compared to alternative solutions, while significantly reducing training time.
+Furthermore, we have identified the optimal number of features necessary to achieve the best possible accuracy.
+Looking ahead, future work could focus on integrating additional interpretability methods with LIME for a more comprehensive feature relevance assessment.
+Extending BOLIMES to multi-omics data, including proteomics and metabolomics, could enhance its applicability.
+Investigating deep learning-based classifiers and validating the approach on larger, diverse datasets and clinical samples would be essential for improving performance and ensuring robustness in real-world applications.
+Acknowledgements: Thanh MA and Thanh-Nghi DO are received support from the European Union’s Horizon research and innovation program under the MSCA-SE (Marie Skłodowska-Curie Actions Staff Exchange) grant agreement 101086252; Call: HORIZON-MSCA-2021-SE-01; Project title: STARWARS.